OpenStudy (zzr0ck3r):

If ai+b = c+d and aj+b = c+e prove that x = (jd-ie)/(d-e) is a solution to ax+b=c

OpenStudy (zzr0ck3r):

this has nothing to do with complex, that is just some variable i.

Directrix (directrix):

What is the title of the course in which you were assigned this problem?

OpenStudy (zzr0ck3r):

um, its from the nine chapters on chinese mathematics, its a 400 level math history class, but this math should not be hard... I just have no idea how to do it for some reason. I try and expand everything and it gets very very very ugly..and I get lost in it.

OpenStudy (zzr0ck3r):

its from chapter 7 in the nine chapters

OpenStudy (zzr0ck3r):

http://www-history.mcs.st-and.ac.uk/HistTopics/Nine_chapters.html scroll down to chapter 7

Directrix (directrix):

When I see i, I think the imaginary number. When I see e I think of another constant. I would have to change the variables to have a chance of working the problem.

OpenStudy (zzr0ck3r):

lol I hear ya, yeah they are all just arbitrary variables.

OpenStudy (zzr0ck3r):

well you know what i mean..

Directrix (directrix):

I do know. Hold up just few seconds while I wrap up two short problems in progress.

OpenStudy (zzr0ck3r):

Ill be here all night:)

Directrix (directrix):

You can move this question to the closed section and type up a new question in a new thread, if there is another question.

OpenStudy (zzr0ck3r):

ah this is all I got, but thanks.

Directrix (directrix):

I am not at all familiar with the "double false position," at least not by that name. There are helpers here who would be.

OpenStudy (zzr0ck3r):

its very strange because my teacher left a note saying just plug in x and varify...lol

OpenStudy (ghazi):

\[x=\frac{ c-b }{ a }\] now find \[\frac{ c-b }{ a } \] from \[ai+b=c+d, ai+b=c+e\]

OpenStudy (ghazi):

sorry aj +b=c+e

OpenStudy (ghazi):

ai-d=c-b and aj-e= c-b now add above equations a(i+j) -(d+e)=2(c-b) now divide both sides by a \[(i+j)-\frac{ d+e }{ a }=2\frac{ c-b }{ a }=2x\] i think now you can do it :)

OpenStudy (ghazi):

also if you subtract ai-d=c-b from aj-e=c-b you will have a(j-i)=e-d therefore\[a=\frac{ e-d }{ j-i }\] now plug in a and rationalise

OpenStudy (ghazi):

and i am sorry if there are sign mistakes :(

OpenStudy (zzr0ck3r):

sweet thnx

OpenStudy (zzr0ck3r):

What am i pluging a into?

OpenStudy (zzr0ck3r):

@ghazi

OpenStudy (zzr0ck3r):

Nm

OpenStudy (ghazi):

well i am sorry i wasnt here , you have to put value of a in to the fraction\[(i+j)-\frac{ d+e }{ a }\] the a in the denominator so that you can rearrange everything and have answer in terms of d and e , i and j only

OpenStudy (zzr0ck3r):

Yeah i got it, tyvm